# Online Hypergraph Matching

### v2

@@ 5 Integral algorithm for bounded degree hypergraphs
$\mathbf{Output:}$ Matching $\mathcal{M}\subset H$
@@
:=$\mathcal{E}:=\Big{\{}\exists v\in h\setminus\{w\}\text{ which is already matched at the arrival of $w$}\Big{\}}.$
@@
    More generally, for every $k\in\{1,\dots,n\}$, the online node $w_{k}$ connects to $n-k+1$ edges $E(w_{k})=E\setminus\{e_{1},\dots,e_{k-1}\}$, and $e_{k}$ is defined as the edge having the lowest fractional value at the end of the iteration of $w_{k}$. We thus get a bound of
    $\sum_{k=1}^{\ell}x(e_{k})\leq\sum_{k=1}^{\ell}\sum_{i=1}^{k}\frac{1}{n-i+1}\qquad\forall\ell\in\{1,\dots,n\}.$ (12)
The inner sum in (12) reaches $\Delta$ approximately when $k\approx(1-e^{-\Delta})n$. For higher values of $k$, it is thus better to use the bound $x(e_{k})\leq\Delta$, which holds by assumption. By defining $p:=\lfloor{e^{-\Delta}n}\rfloor$ and $q:=n-p$, we can now compute a precise upper bound on the total value generated by the algorithm using (12):
$\mathcal{V}(A,\mathcal{H})$ $=\sum_{k=1}^{n}x(e_{k})\leq\sum_{k=1}^{q}\sum_{i=1}^{k}\frac{1}{n-i+1}+\sum_{k=q+1}^{n}\Delta=\sum_{i=1}^{q}\sum_{k=i}^{q}\frac{1}{n-i+1}+p\Delta$
@@
$H_{n}=\ln(n)+\gamma+\epsilon(n)\qquad\text{for some }0<\epsilon(n)<\frac{1}{2n}$ (14)
@@
$H_{n}-H_{p}$ $=\ln\left(\frac{n}{p}\right)+\epsilon(n)-\epsilon(p)\geq\ln\left(\frac{n}{e^{-\Delta}n}\right)-\frac{1}{2p}=\Delta-\frac{1}{2p}$
@@ B.1 Assumption 1: Symmetry
Given a vertex-arrival hypergraph $\mathcal{H}=(V,W,H)$, an automorphism is a permutation of the offline nodes $V$ such that for every $S\subseteq V$ and $w\in W$,
@@
For $S\subseteq V$, we write $\sigma(S)\coloneqq\{\sigma(v):v\in S\}$ for the sake of brevity. For a hyperedge $h=S\cup\{w\}$ where $S\subseteq V$ and $w\in W$, we denote $\sigma(h)\coloneqq\sigma(S)\cup w$. We also denote the relabelled hypergraph after applying $\sigma$ to $\mathcal{H}$ as
$\sigma(\mathcal{H})\coloneqq(\sigma(V),W,\{\sigma(h):h\in H\}).$
@@
Given a vertex-arrival hypergraph $\mathcal{H}=(V,W,H)$, let $\Sigma$ be a subset of its automorphisms. We say that $\mathcal{H}$ is $\Sigma$-symmetric. A fractional matching $x$ in $\mathcal{H}$ is $\Sigma$-symmetric if $x_{h}=x_{\sigma(h)}$ for all $h\in H$ and $\sigma\in\Sigma$. An algorithm $\mathcal{A}$ is $\Sigma$-symmetric on $\mathcal{H}$ if it outputs a $\Sigma$-symmetric fractional matching given $\mathcal{H}$.
Since the construction of our vertex-arrival instance depends on the behaviour of the algorithm, we will overload the notation $\mathcal{H}$ as follows. An (adaptive vertex-arrival) instance is a function $\mathcal{H}$ which takes as input an algorithm $\mathcal{A}$ and outputs a vertex-arrival hypergraph $\mathcal{H}(\mathcal{A})$. For $i\geq 1$, let $\mathcal{H}_{i}(\mathcal{A})$ be the subgraph of $\mathcal{H}(\mathcal{A})$ right after the arrival of the $i$th online node $w_{i}$. We make the following assumptions about $\mathcal{H}$:
  1. 1.
    For any pair of algorithms $\mathcal{A},\mathcal{A}^{\prime}$, we have $\mathcal{H}_{1}(\mathcal{A})=\mathcal{H}_{1}(\mathcal{A}^{\prime})$.
  2. 2.
    For any pair of algorithms $\mathcal{A},\mathcal{A}^{\prime}$ and indices $i,j\geq 1$, we have $V(\mathcal{H}_{i}(\mathcal{A}))=V(\mathcal{H}_{j}(\mathcal{A}^{\prime}))$.
Note that the set of strong automorphisms forms a group.
@@
Let $\Sigma$ be a subgroup of permutations of $V$ such that $\mathcal{H}(\mathcal{A})$ is $\Sigma$-symmetric for every $\Sigma$-symmetric algorithm $\mathcal{A}$. For an algorithm $\mathcal{A}$ on $\mathcal{H}$, let the $\Sigma$-symmetrized algorithm $\operatorname{sym}_{\Sigma}(\mathcal{A})$ be the algorithm which sets:
$x_{h}:=\frac{1}{|\Sigma|}\sum_{\sigma\in\Sigma}x^{\prime}_{\sigma(h)}\qquad\forall h\in H(\mathcal{H})$
where $x^{\prime}_{\sigma(h)}$ is the value that $\mathcal{A}$ would assign to hyperedge $\sigma(h)$ when running on $\mathcal{H}(\operatorname{sym}_{\Sigma}(\mathcal{A}))$.
Observe that $\operatorname{sym}_{\Sigma}(\mathcal{A})$ is $\Sigma$-symmetric, and that $\operatorname{sym}_{\Sigma}(\mathcal{A})$ is a well-defined algorithm for any algorithm $\mathcal{A}$. For $\Sigma$-symmetric algorithms, we have $\operatorname{sym}_{\Sigma}(\mathcal{A})=\mathcal{A}$.
We will now show that one only needs to define the behavior of a $\Sigma$-symmetric instance $\mathcal{H}$ for a $\Sigma$-symmetric algorithm $\mathcal{A}$. That is, any $\Sigma$-symmetric instance $\mathcal{H}$ defined for $\Sigma$-symmetric algorithms can be extended to a $\Sigma$-symmetric instance $\mathcal{H}^{\prime}$ defined for all algorithms.
@@
Let $\mathcal{H}(\mathcal{A})$ be a $\Sigma$-symmetric instance defined for all $\Sigma$-symmetric algorithms $\mathcal{A}$. There is an instance $\mathcal{H}^{\prime}$ defined for all algorithms, such that for any algorithm $\mathcal{A}$:
  • $\mathcal{H}^{\prime}(\mathcal{A})$ is $\Sigma$-symmetric,
@@
We define $\mathcal{H}^{\prime}(\mathcal{A}):=\mathcal{H}(\operatorname{sym}_{\Sigma}(\mathcal{A}))$ for any algorithm $\mathcal{A}$. This immediately gives us the first two properties. This leaves us to show that $\mathcal{V}(\mathcal{A},\mathcal{H}^{\prime})=\mathcal{V}(\operatorname{sym}_{\Sigma}(\mathcal{A}),\mathcal{H})$. Let $x$ be the solution of $\operatorname{sym}_{\Sigma}(\mathcal{A})$ on $\mathcal{H}(\operatorname{sym}_{\Sigma}(\mathcal{A}))$ and let $x^{\prime}$ be the solution of $\mathcal{A}$ on $\mathcal{H}^{\prime}(\mathcal{A})=\mathcal{H}(\operatorname{sym}_{\Sigma}(\mathcal{A}))$. By definition, we have $x_{h}:=\frac{1}{|\Sigma|}\sum_{\sigma\in\Sigma}x^{\prime}_{\sigma(h)}$ for every $h\in H(\mathcal{H})$. So, $\sum_{h\in\mathcal{H}(\operatorname{sym}_{\Sigma}(\mathcal{A}))}x_{h}=\sum_{h\in\mathcal{H}(\operatorname{sym}_{\Sigma}(\mathcal{A}))}\frac{1}{|\Sigma|}\sum_{\sigma\in\Sigma}x^{\prime}_{\sigma(h)}=\sum_{h\in\mathcal{H}(\operatorname{sym}_{\Sigma}(\mathcal{A}))}x^{\prime}_{h}$, which proves the last property. ∎
@@
In Section 4.2, we stated that we assume that the algorithm treats the $k$th vertex in $U_{i}$, say $u_{i,k}$, and the $k$th vertex in $V_{i}$, say $v_{i,k}$, symmetrically. If our constructed hypergraph $\mathcal{H}$ would be symmetric with respect to these vertices, i.e. if the permutation $\sigma$ swapping these two vertices would be a strong automorphism of $\mathcal{H}$, then Lemma B.1 would show that this assumption can be made without loss of generality, as it shows that the instance can be extended so that no algorithm $\mathcal{A}$ can get a better competitive ratio than its symmetric counterpart $\operatorname{sym}(\mathcal{A})$.
@@
However, one part of the instance that breaks this symmetry is the construction given in the proof of Lemma 4.3 and illustrated in Figure 6. As a reminder, this construction is repeatedly applied to submatchings of $\mathcal{M}^{(t)}$ in Section 4.5. Let us fix one such submatching and denote it by $\mathcal{M}:=\mathcal{M}^{(t)}(i,j)$. As described in Section 4.3 and illustrated in Figure 3, if some $u_{i,k}\in\mathcal{M}$, then $v_{i,k}\in\mathcal{M}$ and the submatching is symmetric with respect to this pair, i.e. $\sigma(e)\in\mathcal{M}$ for every edge $e$ in $\mathcal{M}$, where $\sigma$ is the permutation swapping these two vertices. However, due to the Lemma 4.3 construction, $e\cup\{w\}$ might be a hyperedge in $\mathcal{H}$ for some online vertex $w$, while $\sigma(e\cup\{w\})=\sigma(e)\cup\{w\}$ might not be a hyperedge in $\mathcal{H}$.
@@
To fix this, the construction can be slightly tweaked in the following way. An important observation is that the horizontal edges in $\mathcal{M}$ (between $u_{i,k}$ and $v_{i,k}$) are not isomorphic to any other edge in the hypergraph, whereas each of the diagonal edges (all non-horizontal edges) are isomorphic to exactly one other edge in $\mathcal{M}$. For this reason, we can first apply the Lemma 4.3 construction on just the horizontal edges of $\mathcal{M}$.
@@
This results in a constant of $9/2$ instead of $3/2$ inside Lemma 4.3 and a constant of $45$ instead of $15$ inside Lemmas 4.4 and 4.5. This does not affect the asymptotic upper bound for large $m$. Hence, it shows that Theorem 1.1 also holds for non-symmetric algorithms.
@@ B.2 Assumption 2: There is an optimal $\varepsilon$-threshold respecting algorithm
Appendix C Integral upper bound for $k$-uniform hypergraphs

### v3

@@ 5 Integral algorithm for bounded degree hypergraphs
$\mathbf{Output:}$ Matching $\mathcal{M}\subseteq H$
@@
[…]
@@
6 Concluding remarks
Our main contribution was a tight result for the online fractional matching problem under vertex arrivals for $3$-uniform hypergraphs, with matching upper and lower bounds of $(e-1)/(e+1)\approx 0.46$. The biggest open question raised by this work is the integral setting, where the best known algorithm is still the $1/3$-competitive simple greedy algorithm. Possible directions to get an improvement would be to develop new rounding techniques of fractional algorithms on hypergraphs, or by trying to find an analogue of the randomized Ranking algorithm for bipartite graphs. This hypergraph setting adds additional difficulties to the online matching problem, for which new techniques seem to be required. Another interesting direction would be to study this problem in different online arrival models, for instance edge-arrivals.
For $k$-uniform hypergraphs with $k\geq 3$ under vertex arrivals, the integral setting admits upper and lower bounds of $1/k$ and $2/k$, whereas the fractional setting has upper and lower bounds of $\Theta(1/\log k)$. It would still be interesting to derive the optimal closed form solution depending on $k$ for both of these settings, under either vertex or edge arrivals. Note that for large $k$, both models are very similar, as there is a simple reduction from $k$-uniform hypergraphs under edge-arrivals to $(k+1)$-uniform hypergraphs under vertex-arrivals by appending degree one online nodes.
For both the integral and fractional settings, an additional direction would be to study the variant in which previously added hyperedges can be removed from the matching. This is a natural question, as both our upper bound constructions for the fractional and integral settings rely on the fact that an algorithm cannot revoke previous decisions.
@@ 6 Concluding remarks
    More generally, for every $k\in\{1,\dots,n\}$, the online node $w_{k}$ connects to nk+1n-k+1 edges $E(w_{k})=E\setminus\{e_{1},\dots,e_{k-1}\}$, and $e_{k}$ is defined as the edge having the lowest fractional value at the end of the iteration of $w_{k}$. We thus get a bound of
[…]
The inner sum in (11) reaches $\Delta$ approximately when $k\approx(1-e^{-\Delta})n$. For higher values of $k$, it is thus better to use the bound $x(e_{k})\leq\Delta$, which holds by assumption. By defining $p:=\lfloor{e^{-\Delta}n}\rfloor$ and q:=npq:=n-p, we can now compute a precise upper bound on the total value generated by the algorithm using (11):
[…]
@@
[…]
@@
[…]
@@ B.1 Assumption 1: Symmetry
For a vertex-arrival hypergraph $\mathcal{H}=(V,W,H)$, an automorphism is a permutation $\sigma$ of the offline nodes $V$ such that for every $S\subseteq V$ and $w\in W$,
@@
For a subset $S\subseteq V$, we write $\sigma(S)\coloneqq\{\sigma(v):v\in S\}$. For a hyperedge $h=S\cup\{w\}$ where $S\subseteq V$ and $w\in W$, we also write $\sigma(h)\coloneqq\sigma(S)\cup\{w\}$.
@@
For a vertex-arrival hypergraph $\mathcal{H}=(V,W,H)$, let $\Sigma$ be a subset of its automorphisms. We say that a fractional matching $x$ of $\mathcal{H}$ is $\Sigma$-symmetric if $x_{h}=x_{\sigma(h)}$ for all $h\in H$ and $\sigma\in\Sigma$.
Since the construction of our vertex-arrival instance depends on the actions of the algorithm, we will overload the notation $\mathcal{H}$ as follows. A (adaptive) vertex-arrival instance is a function $\mathcal{H}$ which takes as input an algorithm $\mathcal{A}$ and outputs a vertex-arrival hypergraph $\mathcal{H}(\mathcal{A})$. For concreteness, the reader can think of $\mathcal{H}$ as the instance constructed in Section 4. For $i\geq 1$, let $\mathcal{H}_{i}$ be the subinstance of $\mathcal{H}$ that ends with the arrival of the $i$th online node $w_{i}$. We can assume that $\mathcal{H}_{1}(\mathcal{A})=\mathcal{H}_{1}(\mathcal{A}^{\prime})$ for any pair of algorithms $\mathcal{A}$ and $\mathcal{A}^{\prime}$, as algorithms do not take any actions before the first online node arrives. So, the subinstance $\mathcal{H}_{1}$ can be thought of as a hypergraph.
@@
For a vertex-arrival instance $\mathcal{H}$ and an algorithm $\mathcal{A}$, let $\Sigma$ be a subset of automorphisms of $\mathcal{H}(\mathcal{A})$. We say that $\mathcal{A}$ is $\Sigma$-symmetric on $\mathcal{H}$ if it returns a $\Sigma$-symmetric fractional matching of $\mathcal{H}(\mathcal{A})$ given $\mathcal{H}$.
Fix a vertex-arrival instance $\mathcal{H}$, and let $\Sigma$ be a subgroup of automorphisms of $\mathcal{H}_{1}$. For an algorithm $\mathcal{A}$, we define its $\Sigma$-symmetrization $\operatorname{sym}_{\Sigma}(\mathcal{A})$ as follows. We first run $\mathcal{A}$ on $\mathcal{H}_{1}$ to obtain a fractional matching $x^{\prime}$. Then, $\operatorname{sym}_{\Sigma}(\mathcal{A})$ sets
$x_{h}:=\frac{1}{|\Sigma|}\sum_{\sigma\in\Sigma}x^{\prime}_{\sigma(h)}\qquad\forall\;h\in H(\mathcal{H}_{1}).$
This in turn determines the next set of hyperedges in the hypergraph $\mathcal{H}_{2}(\operatorname{sym}_{\Sigma}(\mathcal{A}))$. Note that $\mathcal{H}_{2}(\operatorname{sym}_{\Sigma}(\mathcal{A}))$ may not be equal to $\mathcal{H}_{2}(\mathcal{A})$. As long as $\Sigma$ remains a subgroup of automorphisms of $\mathcal{H}_{2}$, we repeat the process above. So, for any $i\geq 2$, if $x^{\prime}$ is the fractional matching obtained by running $\mathcal{A}$ on $\mathcal{H}_{i}(\operatorname{sym}_{\Sigma}(\mathcal{A}))$, then $\operatorname{sym}_{\Sigma}(\mathcal{A})$ sets
$x_{h}:=\frac{1}{|\Sigma|}\sum_{\sigma\in\Sigma}x^{\prime}_{\sigma(h)}\qquad\forall\;h\in H(\mathcal{H}_{i}(\operatorname{sym}_{\Sigma}(\mathcal{A}))).$
Note that $\operatorname{sym}_{\Sigma}(\mathcal{A})$ is well-defined if and only if $\Sigma$ is a subgroup of automorphisms of the full hypergraph $\mathcal{H}(\operatorname{sym}_{\Sigma}(\mathcal{A}))$.
@@
If $\operatorname{sym}_{\Sigma}(\mathcal{A})$ is well-defined, then it is $\Sigma$-symmetric on $\mathcal{H}$.
Proof.
Let $x$ be the fractional matching obtained by running $\operatorname{sym}_{\Sigma}(\mathcal{A})$ on $\mathcal{H}(\operatorname{sym}_{\Sigma}(\mathcal{A}))$, and let $x^{\prime}$ be the fractional matching obtained by running $\mathcal{A}$ on $\mathcal{H}(\operatorname{sym}_{\Sigma}(\mathcal{A}))$. For any hyperedge $h$ in $\mathcal{H}(\operatorname{sym}_{\Sigma}(\mathcal{A}))$ and any permutation $\tau\in\Sigma$, we have
$x_{h}=\frac{1}{|\Sigma|}\sum_{\sigma\in\Sigma}x^{\prime}_{\sigma(h)}=\frac{1}{|\Sigma|}\sum_{\sigma\in\Sigma}x^{\prime}_{\sigma(\tau(h))}=x_{\tau(h)},$
where the second equality is due to $\Sigma$ being a group. ∎
The following lemma shows that in order to construct a worst-case instance for all algorithms, it suffices to construct a worst-case instance $\mathcal{H}$ for algorithms that are $\Sigma$-symmetric on $\mathcal{H}$. That is, $\mathcal{H}$ can be extended to an instance $\mathcal{H}^{\prime}$ for all algorithms on which every algorithm has the same performance as its $\Sigma$-symmetrization.
Lemma B.2.
For a vertex-arrival instance $\mathcal{H}$, let $\Sigma$ be a subgroup of automorphisms of $\mathcal{H}_{1}$. Suppose that $\Sigma$ remains a subgroup of automorphisms of $\mathcal{H}_{i}(\mathcal{A})$ for any $i\geq 2$ and any algorithm $\mathcal{A}$ that is $\Sigma$-symmetric on $\mathcal{H}_{i-1}$. Then, there exists a vertex-arrival instance $\mathcal{H}^{\prime}$ such that for every algorithm $\mathcal{A}$:
@@
Let $\mathcal{A}$ be any algorithm. Observe that $\operatorname{sym}_{\Sigma}(\mathcal{A})$ is well-defined due to our assumption. Hence, $\operatorname{sym}_{\Sigma}(\mathcal{A})$ is $\Sigma$-symmetric on $\mathcal{H}$ by Lemma B.1,. We define $\mathcal{H}^{\prime}(\mathcal{A}):=\mathcal{H}(\operatorname{sym}_{\Sigma}(\mathcal{A}))$. Let $x$ be the fractional matching obtained by running $\operatorname{sym}_{\Sigma}(\mathcal{A})$ on $\mathcal{H}(\operatorname{sym}_{\Sigma}(\mathcal{A}))$, and let $x^{\prime}$ be the fractional matching obtained by running $\mathcal{A}$ on $\mathcal{H}^{\prime}(\mathcal{A})$. By definition, we have $x_{h}:=\frac{1}{|\Sigma|}\sum_{\sigma\in\Sigma}x^{\prime}_{\sigma(h)}$ for every hyperedge $h\in H(\mathcal{H}(\operatorname{sym}_{\Sigma}(\mathcal{A})))$. Hence,
$\sum_{h}x_{h}=\sum_{h}\frac{1}{|\Sigma|}\sum_{\sigma\in\Sigma}x^{\prime}_{\sigma(h)}=\frac{1}{|\Sigma|}\sum_{\sigma\in\Sigma}\sum_{h}x^{\prime}_{\sigma(h)}=\sum_{h}x^{\prime}_{h}.$
@@
In Section 4.2, we assumed that the algorithm treats the $k$th vertex in $U_{i}$, say $u_{i,k}$, and the $k$th vertex in $V_{i}$, say $v_{i,k}$, symmetrically. If our constructed hypergraph $\mathcal{H}$ was symmetric with respect to these vertices, i.e. if the permutation $\sigma$ swapping $u_{i,k}$ and $v_{i,k}$ for all $i$ and $k$ was an automorphism of $\mathcal{H}$, then Lemma B.2 would show that this assumption can be made without loss of generality. In particular, using the subgroup $\Sigma=(\{\sigma,e\},\circ)$ where $e$ is the identity permutation, it shows that $\mathcal{H}$ can be extended to all algorithms $\mathcal{A}$ so that $\mathcal{A}$ and $\operatorname{sym}_{\Sigma}(\mathcal{A})$ have the same performance.
@@
However, one part of the instance that breaks this symmetry is the construction given in the proof of Lemma 4.4 and illustrated in Figure 7. As a reminder, this construction is repeatedly applied to submatchings of $\mathcal{M}^{(t)}$ in Section 4.5. Let us fix one such submatching and denote it by $\mathcal{M}:=\mathcal{M}^{(t)}(i,j)$. As described in Section 4.3 and illustrated in Figure 4, if some $u_{i,k}\in\mathcal{M}$, then $v_{i,k}\in\mathcal{M}$ and the submatching is symmetric with respect to this pair, i.e. $\sigma(e)\in\mathcal{M}$ for every edge $e$ in $\mathcal{M}$. However, due to the Lemma 4.4 construction, $e\cup\{w\}$ might be a hyperedge in $\mathcal{H}$ for some online vertex $w$, while $\sigma(e\cup\{w\})=\sigma(e)\cup\{w\}$ might not be a hyperedge in $\mathcal{H}$.
@@
To fix this, the construction can be slightly tweaked in the following way. An important observation is that the horizontal edges in $\mathcal{M}$ (between $u_{i,k}$ and $v_{i,k}$) are not isomorphic to any other edge in the hypergraph, whereas each of the diagonal edges (non-horizontal edges) are isomorphic to exactly one other edge in $\mathcal{M}$. For this reason, we can first apply the Lemma 4.4 construction on just the horizontal edges of $\mathcal{M}$.
@@
This results in a constant of $9/2$ instead of $3/2$ in Lemma 4.4, and a constant of $42$ instead of $15$ in Lemmas 4.5, 4.6 and 4.1. This does not affect the asymptotic upper bound for large $m$. Hence, it shows that Theorem 1.1 also holds for non-symmetric algorithms.
@@ B.2 Assumption 2: There is an optimal $\varepsilon$-threshold respecting algorithm
Appendix C Properties of $\psi(t,y)$
Proof of Lemma 4.1.
The first two statement follow directly from the definition and the symmetry of $B(t,\frac{1}{2})$ around $\frac{t}{2}$. For the second statement, let $X\sim B(t,\frac{1}{2})$ and $Y\sim B(1,\frac{1}{2})$ be independent. Then, we have:
$=\sum_{y=0}^{\infty}(y+1)\Pr\left[\left|X-\frac{t}{2}\right|=\frac{y}{2}\right]=1+2\sum_{y=0}^{\infty}\frac{y}{2}\Pr\left[\left|X-\frac{t}{2}\right|=\frac{y}{2}\right]$
$=1+2\>\mathbb{E}\left[\left|X-\frac{t}{2}\right|\right]\leq 1+2\sqrt{\mathbb{E}\left[\left(X-\frac{t}{2}\right)^{2}\right]}\qquad\text{(by Jensen's inequality)}$
$=1+2\sqrt{\operatorname{Var}\left[X\right]}=1+\sqrt{t}.$
Appendix D Integral upper bound for $k$-uniform hypergraphs